If a boat goes downstream 72 miles in 3 hours and upstream 60 miles in 6 hours, the rate of the river and the rate of the boat in still water respectively are? I've got a solution and the answer is correct, but I don't think I worked it out right.
Answer by ankor@dixie-net.com(12701)

If a boat goes downstream 72 miles in 3 hours and upstream 60 miles in 6 hours, the rate of the river and the rate of the boat in still water respectively are?
:
Let x = speed of boat in still water, Let y = speed of river
:
Make two equations; time * speed = distance:
:
Downstream equation:
3(x + y) = 72
Simplify, divide eq by 3
x + y = 24
:
:
Upstream equation:
6(x - y) = 60
:
Simplify divide by 6
x - y = 10
:
Use elimination method on the these two equations:
x + y = 24
x - y = 10
------------ add
2x + 0 = 34
x = 34/2
x = 17 mph, speed of the boat in still water:
:
Find river speed:
17 - y = 10
y = 7 mph is the rive speed
:
:
Check:
3(17+7) = 72
6(17-7) = 60
:
Make sense to you? Any questions?